OpenStudy (anonymous):

find out the answer the integral of cos^10 and show the solution

OpenStudy (anonymous):

\[\int\limits\limits_{}^{} \cos^{10}(x) = 1/2\int\limits\limits_{}^{} (1+\cos(2x))^8 \] You will have to use the half angle formula a whole bunch of times but you can continually simplify it to this form and integrate so you will that have already reduced the power by two

OpenStudy (anonymous):

The solution is pasted above. Check it out and let me know if you have a better method in Half - angle technique as you have indicated above.